As a prospective share holder, you are offered a project that will pay you an income of #8000 in 5years time at 13% rate of discount. What is the present value your future dividend?


1
Expert's answer
2021-08-03T11:52:10-0400

The given information:

Future dividend value: $8,000, time period: 5 years, interest rate: 13% annually

Now the value of dividends at the present period is as follows:

Present value of dividends"=\frac{Future\space value\space of\space dividends}{(1+discount \space rate)^time}"

"=\frac{\$8,000}{(1+0.13)^5}\\=\frac{\$8,000}{(1.13)^5}\\=\frac{\$8,000}{1.8424351793}\\=\$4,342.079488"

Hence the present value of the dividend is $ 4,342.079488